A 1.95m long wire having a mass of 0.100kg is at both ends.The tension in the wire is maintained at 21.0N.




(a).What are the frequencies of the first three allowed modes of vibration?

1
Expert's answer
2021-11-22T10:13:44-0500

The fundamental frequency:


"f_1=\frac{1}{2L}\sqrt{\frac T{m/L}}=2.66\text{ Hz}."

Other two frequencies:


"f_2=2f_1=5.32\text{ Hz},\\
f_3=3f_1=7.98\text{ Hz}."